# Crosstalk-Robust Quantum Control in Multimode Bosonic Systems

### v1

@@ I Introduction
With multiple bosonic modes, the accessible Hilbert space experiences a dramatic expansion. In the context of qudit encoding, the dimension of computational subspace scales as $d^{N}$ for $N$ qudits. For quantum error correction, the integration of multiple bosonic modes facilitates universal quantum computation among protected logical qubits. In most hardware setups designed to realize such multimode systems, a transmon ancilla is directly integrated with various cavity modes [26, 27, 28, 29], as depicted in Fig. 1(a). This architecture allows the ancilla to couple simultaneously with all bosonic modes, thereby facilitating the execution of both single and multimode operations across arbitrary modes. However, this ‘all-to-one’ configuration introduces a significant challenge: crosstalk among the different bosonic modes, as illustrated in Fig. 1(b). Specifically, in the widely-used dispersive regime, the photon population in each cavity mode induces a dispersive shift in the frequency of the transmon ancilla [30]. This becomes a critical issue because standard cavity control protocols depend on driving the ancilla at specific frequencies, which are typically resonant. Consequently, a change in the population of one mode (unknown in a general circuit compilation) can adversely affect the frequency of the ancilla, thereby compromising the fidelity of its controls. This, in turn, can introduce errors in the control of other cavity modes coupled to the ancilla, or even the mode itself. For the purposes of this discussion, we refer to both of these scenarios as crosstalk  [^1] It’s crucial to differentiate this form of crosstalk from that caused by the cross-Kerr effect, which is generally negligible in comparison to the Stark shift-induced crosstalk..
@@
The susceptibility to crosstalk varies according to the nature of the cavity control protocols. The ones that involve a selective ancilla drive with a narrow pulse bandwidth, such as the selective number-dependent arbitrary phase (SNAP) gate [32, 33, 34], are particularly susceptible to crosstalk due to their sensitivity to the ancilla frequency shift. On the other hand, protocols that involve unselective ancilla control, such as the echoed conditional displacement (ECD) [35] and conditional not displacement (CNOD) [36], exhibit reduced sensitivity but still remain susceptible to frequency shifts. In this study, we address the challenge of crosstalk by harnessing the power of quantum optimal control [37, 38, 39, 40]. Our approach involves the development and implementation of robust control, a concept initially introduced in the field of nuclear magnetic resonance [41, 42]. We demonstrate that gates realized by our robust pulse are insensitive to a broad range of frequency variations in the ancilla. We then integrate these robust pulses into a multimode generalization of the ECD protocol, resulting in a notable reduction of crosstalk error, as compared to the standard derivative removal by adiabatic gate (DRAG) pulse [43, 44, 45, 46, 47, 48]. It is crucial to highlight that the use of robust ancilla control is not exclusively applicable to the ECD protocol. Indeed, it can be widely applied to a variety of other cavity control protocols that involve unselective control of the ancilla.
This paper is organized as follows: In Sec. II, we detail the implementation of the multimode ECD protocol, with an emphasis on the spurious phase correction due to cavity-ancilla coupling. Following this, in Sec. III, we employ quantum optimal control to engineer robust ancilla control pulses. Our key findings, which illustrate the suppression of crosstalk, are showcased in Sec. IV. Specifically, we simulate the effect of crosstalk in two representative experimental scenarios: state preparation with spectator modes and a two-mode entanglement operation. We share our conclusions and outlook in Sec. V. Supplementary details are included in the appendices.
@@ II Two-mode ECD protocol with correction of spurious phase accumulation
Universal control of a bosonic mode necessitates an entanglement operation between the linear cavity mode and an additional nonlinear ancilla. One notable example of such an operation is the ECD gate, which was introduced in Ref. 35, 25, and is defined as follows:
$\mathrm{ECD}(\beta)=\hat{D}(\beta/2)\otimes|\text{e}\rangle\langle\text{g}|+\hat{D}(-\beta/2)\otimes|\text{g}\rangle\langle\text{e}|,$ (1)
@@
In practice, to realize a specific target gate or state transfer, we extend the optimization procedure outlined in Ref. 35 to account for the case of two cavity modes. Specifically, we start with a chosen number of unit blocks $N$. The gate sequence consists of $6N$ parameters $\{\beta_{1,n},\beta_{2,n},\varphi_{1,n},\varphi_{2,n},\theta_{1,n},\theta_{2,n}\}$, whose roles are elucidated in Fig. 2(a). Here, the first index in the subscript identifies the cavity mode, while the second index $n\in[1,N]$ refers to the $n$th block in the gate sequence. Subsequently, we optimize these parameters to achieve the requisite circuit-level fidelity. If the optimization fails within a specified number of iterations, we increase the number of unit blocks and repeat the optimization until the desired fidelity is reached.
The above discussion outlines the process to generate the gate sequence for a given task. However, the translation of this abstract sequence into concrete control pulses necessitates an in-depth understanding of the underlying physical systems. In the following, we provide a comprehensive discussion of the driven cavity-ancilla system, and then outline the procedure for constructing the pulse sequence.
@@ II.1 Simulation of the driven coupled cavity-ancilla system
II.1.1 Dispersive Hamiltonian for the undriven system
@@
The above approximations are valid in the regime where $|\Delta_{i}|\gg|K_{\text{q}}|$, consistent with the parameter regime considered in the rest of the paper. Specifically, we fix the parameter set: $K_{\text{q}}/2\pi=-200\,\text{MHz},\Delta_{1}/2\pi=\Delta_{2}/2\pi=2\,\text{GHz}$, and vary the coupling strengths $g_{i}$, taken to be identical for both cavity modes. Given that the dispersive shift $\chi_{i}$ is a directly measurable experimental quantity, we employ it to serve as the proxy for variations in the coupling strengths.
II.1.2 Rotating-frame Hamiltonian for the driven system
[…]
II.1.3 Displaced-frame Hamiltonian for the driven system
When dealing with weak cavity drives, the rotating-frame Hamiltonian can be used to conveniently simulate the system dynamics. However, for large drive amplitudes where the cavity is highly populated, this approach demands a large truncation level of the cavity Hilbert space to ensure numerical convergence, thereby increasing computational cost. A more computationally efficient strategy involves transforming the Hamiltonian to a drive-dependent displaced frame, in which the cavity displacement is eliminated (i.e., terms proportional to $\hat{a}_{i}$ or $\hat{a}_{i}^{\dagger}$ are removed). As a result, simulations in the displaced frame require a considerably smaller dimension for the truncated Hilbert space.
$\partial_{t}\alpha_{i}=i\chi_{i}\alpha_{i}/2+iK_{i}|\alpha_{i}|^{2}\alpha_{i}+iK_{12}\alpha_{i}|\alpha_{\bar{i}}|^{2}-i\Omega_{i}(t),$ (4)
@@ II.1.3 Displaced-frame Hamiltonian for the driven system
[…]
II.1.4 Displaced-frame master equation for the driven open system
$\partial_{t}\eta_{i}=i\chi_{i}\eta_{i}/2+iK_{i}|\eta_{i}|^{2}\eta_{i}+iK_{12}\eta_{i}|\eta_{\bar{i}}|^{2}-i\Omega_{i}(t)-\kappa_{i}\eta_{i}/2,$
@@ II.2 Construction of pulses mitigating spurious phase
Our focus now shifts to the formulation of control pulses realizing universal control of two cavity modes. As illustrated in Fig. 2(a), each unit block (enclosed with dashed lines) comprises paired single-mode ECD gates for every cavity mode, interleaved by ancilla rotations. In this section, we employ standard DRAG pulses for ancilla control, displayed for simplicity as Gaussian shapes in Fig. 2(b). (These pulses can be replaced by the ones obtained from optimal control to improve ancilla gate fidelity, as will be discussed in later sections.) For single-mode ECD gates, we adopt the pulse ansatz introduced in Ref. 35. This ansatz consists of four cavity Gaussian displacement pulses and an ancilla echo pulse in the middle, as depicted in Fig. 2(b). The ansatz is parametrized by the amplitudes $r_{0},r_{1},r_{2},r_{3}$ and the temporal separations $t_{\text{w}}$ of these displacement pulses. Subsequently, we optimize these parameters to generate the desired ECD gates (see details in Ref. 35). This optimization procedure generates control pulses that yield high-fidelity operations for the small cavity nonlinearity parameters taken in Ref. 35.
@@ II.2.1 Characterization of the spurious phase
In scenarios with strong cavity-ancilla coupling and significant cavity displacement, the effects from cavity nonlinearities are non-negligible and need proper treatment. One major consequence of these nonlinearities (self- and cross-Kerrs) is that they result in a time-dependent cavity frequency shift in the form of $\delta_{\text{NL}}(t)\hat{a}^{\dagger}\hat{a}$. This can be inspected in the displaced frame, as shown in $\hat{H}_{\text{diag}}(t)$ in Eq. (6). During the time interval $\tau$, these terms alone lead to relative phase accumulations following $\exp[-i\int_{\tau}\mathrm{d}t^{\prime}n\delta_{\text{NL}}(t^{\prime})]|n\rangle$, where $|n\rangle$ denotes the $n$th Fock state. In the presence of other terms in the displaced Hamiltonian, i.e., $\hat{H}_{\text{static}}$ and $\hat{H}_{\text{off-diag}}(t)$, the above result of phase accumulation is only an approximation, since $\hat{a}^{\dagger}\hat{a}$ terms in $\hat{H}_{\text{diag}}(t)$ do not commute with the rest of the Hamiltonian terms. In the following, we first characterize the approximated phase accumulations from various nonlinearities in a unit block [enclosed by dashed lines in Fig. 2(a)], and then detail the protocol to mitigate them.
@@ II.2.3 Benchmarking the effectiveness of mitigating the spurious phase
Once we implement the phase corrections introduced earlier, the infidelity is shown as the blue solid line. Remarkably, this almost fully eliminates the infidelity arising from spurious phase accumulation due to self-Kerr, cross-Kerr, and second-order dispersive shift. For certain values of $\chi_{i}$, it even marginally outperforms the reference case, thereby validating of our phase-correction strategy.
@@ III Crosstalk error suppression based on quantum optimal control
While the phase correction method detailed in Sec. II significantly reduces infidelity, coherent errors still persist, particularly in cases of strong ancilla-cavity coupling. The primary source of the remaining coherent error in operating multiple cavity modes originates from the crosstalk. This crosstalk is mediated by the coupled transmon ancilla, and arises due to the dispersive shift. Generally, the fidelity of ancilla operations is influenced by the quantum states of the cavity modes. For instance, when a cavity mode is in a Fock state $|n\rangle$, it induces a Stark shift $n\chi$ on the ancilla frequency. Without accounting for this frequency change, the ancilla operations become off-resonant, thereby lowering the fidelity. This error can then propagate through the system, leading to crosstalk between different cavity modes. While integrating tunable couplers into the multimode architecture could potentially mitigate this error [54], our approach instead utilizes resonant drives at the bare ancilla frequency combined with modulation of the pulse envelopes. This approach ensures that the fidelity is insensitive to the detuning between the drive and the Stark-shifted ancilla frequency. Although the concept of Stark shift becomes less definitive when the cavity modes are in superposition or entangled states, robust ancilla control remains a sufficient protocol to mitigate the crosstalk error. In what follows, we first detail our method for optimizing these detuning-robust pulse envelopes for an isolated ancilla. Specifically, we focus on the optimization of $\hat{X}_{\pi/2}$ and $\hat{X}_{\pi}$ gates, which are sufficient for universal control and efficient in experimental calibration (see Appendix B). Subsequently, we integrate these optimized pulses into a composite ancilla-cavity system with minimal modifications, proven effective in reducing crosstalk through numerical simulations.
@@ III.1 Robust control of an isolated transmon ancilla
Detuning-robust control for two-level systems has been extensively investigated with various approaches, such as reverse engineering [55, 56], space curve quantum control [57], pulse engineering [57, 58, 59, 60, 61, 62, 63, 64, 65], and use of the collocation method [66, 67]. However, the protocols developed for two-level systems cannot be directly applied to weakly-anharmonic transmons, primarily due to the leakage out of the qubit subspace. One promising way to address this issue is to impose the DRAG scheme on robust single-quadrature pulses developed for two-level systems [68]. Here, we introduce an optimization-based procedure that provides even better detuning-robust control for transmon $\hat{X}_{\theta}$ gates.

### v2

@@ I Introduction
With multiple bosonic modes, the accessible Hilbert space experiences a dramatic expansion. In the context of qudit encoding, the dimension of computational subspace scales as $d^{N}$ for $N$ qudits. For quantum error correction, the integration of multiple bosonic modes facilitates universal quantum computation among protected logical qubits. In most hardware setups designed to realize such multimode systems, a transmon ancilla is directly integrated with various cavity modes [26, 27, 28, 29], as depicted in Fig. 1(a). This architecture allows the ancilla to couple simultaneously with all bosonic modes, thereby facilitating the execution of both single and multimode operations across arbitrary modes. However, this “all-to-one” configuration introduces a significant challenge: crosstalk among the different bosonic modes, as illustrated in Fig. 1(b). Specifically, in the widely-used dispersive regime, the photon population in each cavity mode induces a dispersive shift in the frequency of the transmon ancilla [30]. This becomes a critical issue because standard cavity control protocols depend on driving the ancilla at specific frequencies, which are typically resonant. Consequently, a change in the population of one mode (unknown in a general circuit compilation) can adversely affect the frequency of the ancilla, thereby compromising the fidelity of its controls. This, in turn, can introduce errors in the control of other cavity modes coupled to the ancilla, or even the mode itself. For the purposes of this discussion, we refer to both of these scenarios as crosstalk  [^1] It’s crucial to differentiate this form of crosstalk from that caused by the cross-Kerr effect, which is generally negligible in comparison to the Stark shift-induced crosstalk..
@@
The susceptibility to crosstalk varies according to the nature of the cavity control protocols. The ones that involve a selective ancilla drive with a narrow pulse bandwidth, such as the selective number-dependent arbitrary phase (SNAP) gate [32, 33, 34], are particularly susceptible to crosstalk. This susceptibility arises from their sensitivity to the ancilla frequency shift, even with the help of optimal control [35]. On the other hand, protocols that involve unselective ancilla control, such as the echoed conditional displacement (ECD) [36] and conditional not displacement (CNOD) [37], exhibit reduced sensitivity but still remain susceptible to frequency shifts. In this study, we address the challenge of crosstalk by harnessing the power of quantum optimal control [38, 39, 40, 41]. Our approach involves the development and implementation of robust control, a concept initially introduced in the field of nuclear magnetic resonance [42, 43]. We demonstrate that gates realized by our robust pulse are insensitive to a broad range of frequency variations in the ancilla. We then integrate these robust pulses into a multimode generalization of the ECD protocol, resulting in a notable reduction of crosstalk error, as compared to the standard derivative removal by adiabatic gate (DRAG) pulse [44, 45, 46, 47, 48, 49]. It is crucial to highlight that the use of robust ancilla control is not exclusively applicable to the ECD protocol. Indeed, it can be widely applied to a variety of other cavity control protocols that involve unselective control of the ancilla.
This paper is organized as follows: In Sec. II, we detail the implementation of the multimode ECD protocol, with an emphasis on the spurious phase correction due to cavity-ancilla coupling. Following this, in Sec. III, we employ quantum optimal control to engineer robust ancilla control pulses. Our key findings, which illustrate the suppression of crosstalk, are showcased in Sec. IV. Specifically, we simulate the effect of crosstalk in two representative experimental scenarios: state preparation with spectator modes and a two-mode entanglement operation. We discuss and compare our multimode architecture with other existing ones in the literature in Sec. V, before sharing our conclusions in Sec. VI. Supplementary details are included in the appendices.
@@ II Two-mode ECD protocol with correction of spurious phase accumulation
Universal control of a bosonic mode necessitates an entanglement operation between the linear cavity mode and an additional nonlinear ancilla. One notable example of such an operation is the ECD gate, which was introduced in Ref. 36, 25, and is defined as follows:
$\hat{D}(\beta/2)\otimes|\text{e}\rangle\langle\text{g}|+\hat{D}(-\beta/2)\otimes|\text{g}\rangle\langle\text{e}|,$ (1)
@@
In practice, to realize a specific target gate or state transfer, we generalize the optimization procedure outlined in Ref. 36 for a single mode to account for the case of two cavity modes. Specifically, we start with a chosen number of unit blocks $N$. The gate sequence consists of $6N$ parameters $\{\beta_{1,n},\beta_{2,n},\varphi_{1,n},\varphi_{2,n},\theta_{1,n},\theta_{2,n}\}$, whose roles are elucidated in Fig. 2(a). Here, the first index in the subscript identifies the cavity mode, while the second index $n\in[1,N]$ refers to the $n$th block in the gate sequence. Subsequently, we optimize these parameters to achieve the requisite circuit-level fidelity. If the optimization fails within a specified number of iterations, we increase the number of unit blocks and repeat the optimization until the desired fidelity is reached. The numerical implementation is similar to the case for a single cavity mode. See Appendix S7 in Ref. 36 for more details.
The above discussion outlines the process of generating the gate sequence for a given task. However, translating this abstract sequence into concrete control pulses requires a thorough understanding of the underlying physical systems. In the following, we provide a brief overview of simulating a driven cavity-ancilla system (with detailed derivations in Appendix B), and then outline the procedure for constructing the corresponding pulse sequence.
@@ II.1 Simulation of the driven coupled cavity-ancilla system
Within the dispersive regime, the coupled cavity-ancilla system can be well described by the following Hamiltonian:
@@
$+K_{12}\hat{a}_{1}^{\dagger}\hat{a}_{1}\hat{a}_{2}^{\dagger}\hat{a}_{2}+\omega_{\text{q}}\hat{q}^{\dagger}\hat{q}+\frac{K_{\text{q}}}{2}\hat{q}^{{\dagger}2}\hat{q}^{2},$
where $\omega_{\text{q}}$ and $\omega_{\text{i}}$ denote the frequencies of the ancilla and cavity modes. The nonlinearities are defined as follows. The self- and cross-Kerr interaction strengths for the two cavity modes are described by $K_{i}$ and $K_{12}$, and the first- and second-order dispersive shifts between the cavity modes and the ancilla are symbolized by $\chi_{i}$ and $\chi_{i}^{\prime}$, respectively. Note that the above nonlinearities are inherited from the transmon anharmonicity, and are interconnected through the relations in Eq. (26). The effects of higher-order transmon nonlinearities are analyzed in Appendix C. Although our protocol does not necessitate identical cavity parameters, we take them to be identical for simplicity in the numerical simulations throughout this paper.
To execute ECD gates, microwave drives are applied on both the ancilla and the two cavity modes. In particular, the ancilla is driven resonantly with temporal profile $\varepsilon(t)$, while the two cavity modes are driven off-resonantly with detunings $\omega_{i}-\omega_{\text{d},i}=-\chi_{i}/2$ and temporal profiles $\Omega_{i}(t)$. To simulate the driven dynamics efficiently, it is advantageous to first go into the rotating frame defined by both the cavity and ancilla drive frequencies, and then the displaced frame, characterized by the unitary transformation $\hat{U}_{\alpha}(t)=\Pi_{i=1,2}\hat{D}_{i}^{\dagger}[\alpha_{i}(t)]$. Here, the displacement $\alpha_{i}(t)$ is obtained by solving the trajectory equation in Eq. (28). The resulting displaced frame Hamiltonian is detailed in Appendix B.3. In addition to those static terms, including the dispersive interaction $\chi_{i}\hat{a}_{i}^{\dagger}\hat{a}_{i}$ leading to the conditional displacement in Eq. (1), there exists time-dependent diagonal terms, which depend on the displacement $\alpha_{i}(t)$:
@@
In the presence of strong cavity nonlinearities, the above terms introduce significant corrections to the system dynamics, necessitating careful consideration in the design of high-fidelity cavity control pulses.
@@ II.2 Construction of pulses mitigating spurious phase
Our focus now shifts to the formulation of control pulses realizing universal control of two cavity modes. As illustrated in Fig. 2(a), each unit block (enclosed with dashed lines) comprises paired single-mode ECD gates for every cavity mode, interleaved by ancilla rotations. In this section, we employ standard DRAG pulses for ancilla control, displayed for simplicity as Gaussian shapes in Fig. 2(b). (These pulses can be replaced by the ones obtained from optimal control to improve ancilla gate fidelity, as will be discussed in later sections.) For single-mode ECD gates, we adopt the pulse ansatz introduced in Ref. 36. This ansatz consists of four cavity Gaussian displacement pulses and an ancilla echo pulse in the middle, as depicted in Fig. 2(b). The ansatz is parametrized by the amplitudes $r_{0},r_{1},r_{2},r_{3}$ and the temporal separations $t_{\text{w}}$ of these displacement pulses. We then follow the same routine in Ref. 36 to optimize these parameters for desired ECD gates. This optimization procedure generates control pulses that yield high-fidelity operations for the small cavity nonlinearity parameters taken in Ref. 36.
@@ II.2.1 Characterization of the spurious phase
In scenarios with strong cavity-ancilla coupling and significant cavity displacement, the effects from cavity nonlinearities are non-negligible and need proper treatment. One major consequence of these nonlinearities (self- and cross-Kerrs) is that they result in a time-dependent cavity frequency shift in the form of $\delta_{\text{NL}}(t)\hat{a}_{i}^{\dagger}\hat{a}_{i}$. This can be inspected in the displaced frame, as shown in $\hat{H}_{\text{diag}}(t)$ in Eq. (2). During the time interval $\tau$, these terms alone lead to relative phase accumulations following $\exp[-i\int_{\tau}\mathrm{d}t^{\prime}n_{i}\delta_{\text{NL}}(t^{\prime})]|n_{i}\rangle$, where $|n_{i}\rangle$ denotes the $n$th Fock state of the $i$th cavity mode. However, when off-diagonal terms are present [see Eq. (30)], describing the spurious dynamics solely by phase accumulation is an approximation. These off-diagonal terms induce additional spurious dynamics, such as ancilla-state dependent displacement (e.g., $\hat{a}_{i}{\hat{q}}^{\dagger}\hat{q}$) and squeezing (e.g., $\hat{a}_{i}^{2}\hat{q}^{\dagger}\hat{q}$), which do not commute with the diagonal terms. Together, these effects make the phase accumulation description approximate. In the following, we first characterize the approximated phase accumulations from various nonlinearities in a unit block [enclosed by dashed lines in Fig. 2(a)], and then detail the protocol to mitigate them.
@@ II.2.3 Benchmarking the effectiveness of mitigating the spurious phase
Once we implement the phase corrections introduced earlier, the infidelity is shown as the blue solid line. Remarkably, this almost fully eliminates the infidelity arising from spurious phase accumulation due to self-Kerr, cross-Kerr, and second-order dispersive shift. For certain values of $\chi_{i}$, it even marginally outperforms the reference case, thereby validating our phase-correction strategy.
@@ III Crosstalk error suppression based on quantum optimal control
While the phase correction method detailed in Sec. II significantly reduces infidelity, coherent errors still persist, particularly in cases of strong ancilla-cavity coupling. The primary source of the remaining coherent error in operating multiple cavity modes originates from the crosstalk. This crosstalk is mediated by the coupled transmon ancilla, and arises due to the dispersive shift. Generally, the fidelity of ancilla operations is influenced by the quantum states of the cavity modes. For instance, when a cavity mode is in a Fock state $|n\rangle$, it induces a Stark shift $n\chi$ on the ancilla frequency. Without accounting for this frequency change, the ancilla operations become off-resonant, thereby lowering the fidelity. This error can then propagate through the system, leading to crosstalk between different cavity modes. While integrating tunable couplers into the multimode architecture could potentially mitigate this error [53], our approach instead utilizes resonant drives at the bare ancilla frequency combined with modulation of the pulse envelopes. This approach ensures that the fidelity is insensitive to the detuning between the drive and the Stark-shifted ancilla frequency. Although the concept of Stark shift becomes less definitive when the cavity modes are in superposition or entangled states, robust ancilla control remains a sufficient protocol to mitigate the crosstalk error. In what follows, we first detail our method for optimizing these detuning-robust pulse envelopes for an isolated ancilla. Specifically, we focus on the optimization of $\hat{X}_{\pi/2}$ and $\hat{X}_{\pi}$ gates, which are sufficient for universal control and efficient in experimental calibration (see Appendix D). Subsequently, we integrate these optimized pulses into a composite ancilla-cavity system with minimal modifications, proven effective in reducing crosstalk through numerical simulations.
@@ III.1 Robust control of an isolated transmon ancilla
Detuning-robust control for two-level systems has been extensively investigated using various approaches, including reverse engineering [54, 55], space curve quantum control [56], pulse engineering [56, 57, 58, 59, 60, 61, 62, 63, 64], and the collocation method [65, 66]. However, protocols developed for two-level systems cannot be directly applied to weakly-anharmonic transmons due to leakage out of the qubit subspace. One promising solution is to apply the DRAG scheme to robust single-quadrature pulses developed for two-level systems [67]. Additionally, Ref. [68] directly optimizes two-quadrature pulses for transmons. In this work, we introduce an alternative optimization-based procedure to design detuning-robust control for transmon $\hat{X}_{\theta}$ gates and benchmark its performance. A detailed numerical comparison with selected literature is presented in Appendix E.